The lenth of a rug is 6ft more than its width. the area is 40sq ft. Find the dimensions of the rug.

Answer by mananth(16946)

let width be w
length = w+6
...
area = L*W
w(w+6)=40
w^2+6x=40
w^2+6x-40=0
(w+10)(w-4)=0
w=4 the width
length = 10 ft
..
m.ananth@hotmail.ca